### v1

# Multi-Agent Inverted Transformer for Flight Trajectory Prediction

Abstract
Flight trajectory prediction for multiple aircraft is essential and provides critical insights into how aircraft navigate within current air traffic flows. However, predicting multi-agent flight trajectories is inherently challenging. One of the major difficulties is modeling both the individual aircraft behaviors over time and the complex interactions between flights. Generating explainable prediction outcomes is also a challenge. Therefore, we propose a Multi-Agent Inverted Transformer, MAIFormer, as a novel neural architecture that predicts multi-agent flight trajectories. The proposed framework features two key attention modules: (i) masked multivariate attention, which captures spatio-temporal patterns of individual aircraft, and (ii) agent attention, which models the social patterns among multiple agents in complex air traffic scenes. We evaluated MAIFormer using a real-world automatic dependent surveillance-broadcast flight trajectory dataset from the terminal airspace of Incheon International Airport in South Korea. The experimental results show that MAIFormer achieves the best performance across multiple metrics and outperforms other methods. In addition, MAIFormer produces prediction outcomes that are interpretable from a human perspective, which improves both the transparency of the model and its practical utility in air traffic control.
I Introduction
The growing demand for air travel has resulted in greater airspace congestion, which necessitates advanced air traffic management (ATM) strategies, such as trajectory-based operation (TBO) [1]. In TBO, aircraft operations are managed based on their trajectories in four-dimensional spaces (latitude, longitude, altitude, and time), which provides more precise and reliable trajectory information that air traffic controllers (ATCs) can use to optimize their decisions. In this context, accurate and reliable flight trajectory prediction is increasingly important.
Conventionally, physics-based methods for flight trajectory prediction have been widely used, which rely on kinetic models of aircraft and aerodynamic performance data such as the Base of Aircraft Data (BADA) [2, 3, 4, 5]. These approaches are effective in low-density airspace under normal conditions, but they often struggle to predict flight trajectories accurately when they deviate substantially from standard flight procedures due to interventions by ATCs [6]. Therefore, various data-driven approaches have gained significant attention in recent years.
Early models based on classical multivariate regression techniques have been proposed for flight trajectory prediction [7, 8]. Several studies have explored the use of Gaussian mixture models (GMMs) to represent the statistical distributions of flight trajectories and apply them to flight trajectory prediction [9, 10, 11, 12, 13].
Recent advancements have increasingly focused on deep learning-based approaches. Several models for flight trajectory prediction based on long short-term memory (LSTM) networks have been proposed [14, 15, 16]. Other studies have combined convolutional neural networks (CNNs) with LSTMs to improve the extraction of the spatio-temporal patterns of flight trajectories [17, 18].
More recently, Transformer-based models have gained prominence due to their better capabilities in sequence representation. Vanilla Transformer architectures have been employed for long-term flight trajectory prediction [19], and the temporal fusion Transformer has been used to improve the capture of temporal dependencies in flight trajectories [20]. Other approaches have explored alternative representations of flight trajectories by applying binary encoding or inverted embedding techniques [21, 22]. The agent-aware attention mechanism originally introduced in AgentFormer [23] has been adapted for multi-agent flight trajectory prediction [24, 25].
Building upon previous studies, this paper proposes a Multi-Agent Inverted Transformer (MAIFomer) for flight trajectory prediction. Aircraft movements in airspace are inherently multi-agent in nature, and an effective flight trajectory prediction model must account for both the spatio-temporal dynamics of individual aircraft and their social interactions with surrounding aircraft. The standard attention-based approach to modeling the interactions typically relies on computing attention scores between trajectory points in different time steps across different aircraft [24, 25]. Although this approach has demonstrated competitive performance in improving the prediction accuracy, it often yields attention distributions that are either irrelevant or difficult to interpret from a human perspective. In contrast, the proposed MAIFormer treats the entire trajectory of each aircraft as a single token and computes attention between agents (i.e., aircraft) to capture their interactions, as illustrated in Figure 1.
Figure 1: Conventional approaches model inter-agent interactions by applying attention mechanisms across trajectory points, whereas MAIFormer models these interactions through direct agent-to-agent attention.
The MAIFormer architecture has two distinct attention stages. In the first stage, masked multivariate attention (MMA) is used to capture the series-wise relationships among variates for each individual aircraft trajectory. In the second stage, agent attention (AA) is applied across different flights to model social interactions between aircraft. Unlike conventional approaches, MAIFormer decouples the processes of modeling intra-agent and inter-agent relationships into separate attention stages. This separation improves the accuracy of the prediction outputs and the interpretability of attention distributions. Interpretability is particularly important in safety-critical domains such as ATM, where the outputs of a flight trajectory prediction model must be understandable by human operators such as ATCs.
The main contributions of this work are the following:
  • We propose MAIFormer, a novel multi-agent aircraft trajectory model that incorporates two specialized attention modules to reflect the hierarchical structure of multi-agent flight interactions.
  • The proposed model demonstrates significant performance improvements in prediction accuracy over other methods when evaluated on a real-world automatic dependent surveillance-broadcast (ADS-B) flight trajectory dataset, including the single-aircraft Transformer and AgentFormer.
  • MAIFormer uses a hierarchical attention mechanism to assign a single interpretable attention score to each aircraft, which enables a more intuitive understanding of how the model perceives inter-agent interactions within a traffic scene. This abstraction improves the model’s transparency and usability for human operators, such as ATCs.
II Methods
II-A Problem Formulation
Let $\mathbf{X}$ represent the input air traffic scene involving $N$ aircraft over a time horizon $T$. The state of the $n$-th aircraft in the air traffic scene at time $t$ is given by a feature vector $\mathbf{x}_{n}^{t}\in\mathbb{R}^{1\times F}$, where $F$ is the number of variates (e.g., latitude, longitude, altitude). The past trajectory sequence of the $n$-th aircraft over a time horizon $T$ is denoted by $\mathbf{X}_{n}=(\mathbf{x}_{n}^{1},\mathbf{x}_{n}^{2},\dots,\mathbf{x}_{n}^{T})\in\mathbb{R}^{T\times F}$. Similarly, the future trajectory sequence of the $n$-th aircraft over a prediction horizon $S$ is denoted by $\mathbf{Y}_{n}=(\mathbf{y}_{n}^{1},\mathbf{y}_{n}^{2},\dots,\mathbf{y}_{n}^{S})\in\mathbb{R}^{S\times F}$. Therefore, the full multi-agent past and future air traffic scenes are represented as $\mathbf{X}=(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{N})\in\mathbb{R}^{N\times T\times F}$ and $\mathbf{Y}=(\mathbf{Y}_{1},\mathbf{Y}_{2},\dots,\mathbf{Y}_{N})\in\mathbb{R}^{N\times S\times F}$, respectively. Our goal is first to train a multi-agent flight trajectory prediction model $\mathcal{F}_{\theta}(\mathbf{Y}|\mathbf{X})$, which minimizes the difference between $\mathbf{Y}$ and $\widehat{\mathbf{Y}}$, where $\mathcal{F}$ is the multi-agent flight trajectory prediction model, $\theta$ represents the learnable model parameters, and $\mathbf{Y}$ and $\widehat{\mathbf{Y}}$ are the actual and predicted trajectories of multi-agent aircraft, respectively. Formally, the overall learning problem is defined as:
$\theta^{*}=\min_{\theta}\mathcal{L}(\mathbf{Y},\widehat{\mathbf{Y}}),~~~{\rm s.t.}~~~\widehat{\mathbf{Y}}\sim\mathcal{F}_{\theta}.$ (1)
where $\mathcal{L}$ is a loss function that quantifies the difference between $\mathbf{Y}$ and $\widehat{\mathbf{Y}}$.
Figure 2: Overall architecture of MAIFormer. The past air traffic scene is first tokenized into variate tokens. The sequence of variate tokens from all agents is fed into stacked MAIFormer layers where spatio-temporal and social interaction patterns of the air traffic scene are captured with masked multivariate attention and agent attention, respectively.
II-B Overview of the Proposed Framework
Figure 2 illustrates the overall architecture of the proposed method, which has three main components. First, a novel tokenization module embeds air traffic scene $\mathbf{X}$ into a sequence of variate tokens using an inverted embedding strategy [26]. Second, MAIFormer, an encoder-only Transformer, extracts both the intra-agent and inter-agent relationships from the multi-agent flight trajectories within the air traffic scene. Finally, given the latent representation of $\mathbf{X}$ produced by the MAIFormer, a multi-layer perceptron (MLP) decoder is used to predict the future air traffic scene $\widehat{\mathbf{Y}}$.
MAIFormer is a fully Transformer-based encoder, and its core operation is centered around self attention (SA) [27]. SA computes pairwise relationships between elements in an input sequence to determine how much attention each element should receive relative to the others. Mathematically, SA is defined as:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{QK}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}$ (2)
where $\mathbf{Q}\in\mathbb{R}^{T\times D},\mathbf{K}\in\mathbb{R}^{T\times D}$, and $\mathbf{V}\in\mathbb{R}^{T\times D}$ represent queries, keys, and values, respectively. $D$ is the model’s latent dimension, and $d_{k}$ is the dimensionality of the keys.
II-C Air Traffic Scene Tokenization
Given the multi-agent flight trajectory sequences $\mathbf{X}\in\mathbb{R}^{N\times T\times F}$, we invert and reshape it into $\mathbf{\overline{X}}\in\mathbb{R}^{(N\cdot F)\times T}$ and treat the trajectories from the multiple aircraft as a single multivariate time-series with $N\cdot F$ variates. We then embed $\mathbf{\overline{X}}$ along the time dimension as follows:
$\mathbf{C}=\text{SceneTokenization}(\mathbf{\overline{X}})$ (3)
$\mathbf{C}=(\mathbf{c}^{1}_{1},\mathbf{c}^{2}_{1},\dots,\mathbf{c}^{F}_{1},\dots,\mathbf{c}^{1}_{N},\dots\mathbf{c}^{F}_{N})\in\mathbb{R}^{(N\cdot F)\times D}$ denotes the sequence of variate tokens, where each token is generated by applying a linear embedding to the original time series of each variate in $\mathbf{\overline{X}}$. Notably, the positional encoding is not applied after the tokenization, since there is no inherent ordering among agents in the air traffic scene.
Figure 3: Air traffic scene tokenization. We apply an inverted embedding to each agent’s past trajectory and generate a sequence of variate tokens from all agents. In this example, both the number of agents NN and the number of variates $F$ are set to three, which results in three tokens per agent and a total of nine tokens after tokenization. For consistency, we maintain these settings (three agents and three variates) in the subsequent figures.
II-D MAIFormer: Multi-Agent Inverted Transformer
MAIFormer consists of $L$ identical layers, where each layer comprises three key components: an MMA module, an AA module, and a feed-forward network (FFN) module. We define each MAIFormer layer as follows:
$\mathbf{C}_{\text{ST}}^{\ell-1}$ $=\text{MaskedMultivariateAttention}(\text{LN}(\mathbf{C}^{\ell-1})),$ (4)
$\mathbf{C}_{\text{SC}}^{\ell-1}$ $=\text{AgentAttention}(\text{LN}(\mathbf{C}_{\text{A}}^{\ell-1}))+\mathbf{C}_{\text{A}}^{\ell-1},$ (5)
$\mathbf{C}^{\ell}$ $=\text{FFN}(\text{LN}(\mathbf{C}_{\text{SC}}^{\ell-1}))+\mathbf{C}_{\text{SC}}^{\ell-1},\quad\ell=(1,\ldots,L)$ (6)
where $\mathbf{C}^{\ell}$ is the sequence of variate tokens from the $\ell$-th layer, and the output of the MMA module is $\mathbf{C}_{\text{ST}}$. The sequence of agent tokens $\mathbf{C}_{\text{A}}$ is then obtained by reshaping $\mathbf{C}_{\text{ST}}$ through agent-wise concatenation, and the output of the AA module is $\mathbf{C}_{\text{SC}}$. To improve the convergence and training stability, layer normalization (LN) was applied to the input of each module.
The FFN in MAIFormer was designed with shared weights across all tokens, which allows the model to learn intrinsic temporal patterns across different variates. The FFN operation is defined as:
$\text{FFN}(\mathbf{C}^{\ell-1}_{\text{SC}})=\text{GELU}(\mathbf{C}^{\ell-1}_{\text{SC}}\mathbf{W}^{1}+\boldsymbol{b}^{1})\mathbf{W}^{2}+\boldsymbol{b}^{2}$ (7)
where $\mathbf{W}^{1}\in\mathbb{R}^{D\times(4\cdot D)}$ and $\mathbf{W}^{2}\in\mathbb{R}^{(4\cdot D)\times D}$ are learnable projection matrices, while $\boldsymbol{b}^{1}\in\mathbb{R}^{4\cdot D}$ and $\boldsymbol{b}^{2}\in\mathbb{R}^{D}$ are the corresponding bias vectors. Gaussian error linear unit (GELU) [28] is used as the nonlinear activation function in the FFN and defined as follows:
$\text{GELU}(x)=x\cdot\Phi(x)$ (8)
where $\Phi(x)$ is the cumulative distribution function (CDF) of the standard normal distribution. The pseudo-code of MAIFormer is presented in Algorithm 1.
Algorithm 1 Pseudo-code of MAIFormer
0: Tokenized air traffic scene $\mathbf{C}\in\mathbb{R}^{(N\cdot F)\times D}$
1:for $l\leftarrow 1$ to $L$ do
2:  $\mathbf{C}_{\text{ST}}^{\ell-1}=\text{MMA}(\text{LN}(\mathbf{C}^{\ell-1}))$ $\triangleright\mathbf{C}_{\text{ST}}^{\ell-1}\in\mathbb{R}^{(N\cdot F)\times D}$
3:  $\mathbf{C}_{\text{A}}^{\ell-1}=\mathbf{C}_{\text{ST}}^{\ell-1}.\text{reshape}$ $\triangleright\mathbf{C}_{\text{A}}^{\ell-1}\in\mathbb{R}^{N\times(F\cdot D)}$
4:  $\mathbf{C}_{\text{SC}}^{\ell-1}=\text{AA}(\text{LN}(\mathbf{C}_{\text{A}}^{\ell-1}))+\mathbf{C}_{\text{A}}^{\ell-1}$ $\triangleright\mathbf{C}_{\text{SC}}^{\ell-1}\in\mathbb{R}^{N\times(F\cdot D)}$
5:  $\mathbf{C}_{\text{SC}}^{\ell-1}=\mathbf{C}_{\text{SC}}^{\ell-1}.\text{reshape}$ $\triangleright\mathbf{C}_{\text{SC}}^{\ell-1}\in\mathbb{R}^{(N\cdot F)\times D}$
6:  $\mathbf{C}^{\ell}=\text{FFN}(\text{LN}(\mathbf{C}_{\text{SC}}^{\ell-1}))+\mathbf{C}_{\text{SC}}^{\ell-1}$ $\triangleright\mathbf{C}^{\ell}\in\mathbb{R}^{(N\cdot F)\times D}$
7:end for
II-D1 Masked Multivariate Attention
MMA was specifically designed to model the spatio-temporal (ST) patterns within each individual flight trajectory. In MMA, when computing attention scores between variate tokens for the $i$-th agent $\mathbf{C}_{i}=(\mathbf{c}^{1}_{i},\mathbf{c}^{2}_{i},\dots,\mathbf{c}^{F}_{i})\in\mathbb{R}^{F\times D}$, the variate tokens of other agents are masked using a mask matrix $\mathbf{M}\in\mathbb{R}^{(N\cdot F)\times(N\cdot F)}$, which is defined as:
[…]
where $\left\lfloor\cdot\right\rfloor$ denotes the floor function. Using the mask matrix $\mathbf{M}$, we can ensure that the attentions in MMA are only computed among variates within the same aircraft, and the attentions to the variates from other aircraft are restricted. Since the mixing of variate information between different aircraft is prevented in MMA, MAIFormer more effectively captures fine-grained individual aircraft behavior, which improves prediction accuracy. The overall MMA process is illustrated in Figure 4 (left).
Figure 4: Illustration of masked multivariate attention (MMA) (left) and agent attention (AA) (right). In the MMA, the mask matrix $\mathbf{M}$ restricts the attention scores only to the variates from the same agent, which prevents information sharing across agents. In contrast, the AA allows all agents to share information with one another.
II-D2 Agent Attention
AA was designed to model the social context (SC) patterns between agents. To this end, SA is applied across agent tokens, which are constructed by agent-wise concatenation of the variate tokens generated from the previous MMA stage. The agent token of the $i$-th agent is defined as:
$\mathbf{C}_{\text{A},i}=(\mathbf{c}_{\text{ST},i}^{1}\oplus\mathbf{c}_{\text{ST},i}^{2}\oplus\cdots\oplus\ \mathbf{c}_{\text{ST},i}^{F})\in\mathbb{R}^{1\times(F\cdot D)}$ (10)
where $\mathbf{c}_{\text{ST},i}^{j}$ represents the $j$-th variate token of the $i$-th agent, and $\oplus$ denotes concatenation along the variate dimension. Once we obtain a whole sequence of agent tokens $\mathbf{C}_{\text{A}}\in\mathbb{R}^{N\times(F\cdot D)}$ for $N$ agents, AA is applied to that sequence as follows:
$\mathbf{C}_{\text{SC}}=\text{AgentAttention}(\mathbf{Q}_{\text{SC}},\mathbf{K}_{\text{SC}},\mathbf{V}_{\text{SC}})$ (11)
$\mathbf{Q}_{\text{SC}},\mathbf{K}_{\text{SC}},\mathbf{V}_{\text{SC}}=\mathbf{C}_{\text{A}}\mathbf{W}^{\mathbf{Q}}_{\text{SC}},\mathbf{C}_{\text{A}}\mathbf{W}^{\mathbf{K}}_{\text{SC}},\mathbf{C}_{\text{A}}\mathbf{W}^{\mathbf{V}}_{\text{SC}}$
where each $\mathbf{W}_{\text{SC}}\in\mathbb{R}^{(F\cdot D)\times(F\cdot D)}$ is a learnable projection for agent tokens that transforms the agent tokens into agent queries $\mathbf{Q}_{\text{SC}}$, agent keys $\mathbf{K}_{\text{SC}}$, and agent values $\mathbf{V}_{\text{SC}}$.
The agent-level representations in AA allows inter-agent relationships to be effectively captured without the complexity of trajectory point-level attentions across agents, which often obscure attention interpretability. The output vector from AA is then reshaped back to its original shape, $\mathbf{C}_{\text{SC}}\in\mathbb{R}^{(N\cdot F)\times D}$, to be used in the subsequent operations in the next FFN module. The overall process of AA is illustrated in Figure 4 (right).
II-D3 MLP Decoder
The latent representations $\mathbf{C}^{L}\in\mathbb{R}^{(N\cdot F)\times D}$ from the last MAIFormer layer are passed through a MLP decoder for prediction:
$\widehat{\mathbf{Y}}=\text{MLP}(\mathbf{C}^{L})$ (12)
The MLP consists of multiple linear layers with GELUs, which progressively reduce the dimension from $\mathbb{R}^{D}$ to $\mathbb{R}^{S}$. The output $\widehat{\mathbf{Y}}\in\mathbb{R}^{(N\cdot F)\times S}$ is then reshaped and transposed to $\mathbb{R}^{N\times S\times F}$ to match the original shape of the multi-agent trajectories. The MLP decoder enables simultaneous non-autoregressive generation of trajectories for all aircraft, which effectively mitigates the error accumulation that is typically observed in autoregressive models.
III Experimental Setup
III-A Dataset Description
We evaluated the performance of the proposed method using ADS-B flight trajectory dataset obtained from FlightRadar24111https://www.flightradar24.com/. The dataset spatially covers the terminal airspace within a radius of 70 nautical miles (NM) of Incheon International Airport (ICN) in South Korea from January 2023 to May 2023 (five months). Each flight trajectory data includes a timestamp, aircraft type, a sequence of geographical positions (latitude, longitude, and altitude), and other operationally relevant information such as the departure and arrival airports. In this work, we focus exclusively on arrival flights as arrival operations at ICN are considerably more complex than departures, which makes multi-agent interactions more critical. Notably, arrival and departure traffic flows are spatially segregated via dedicated flight procedures and are typically managed by different ATCs, which minimizes their interaction. Therefore, restricting the dataset to arrivals does not compromise the validity of our analysis.
To address the irregular sampling rate of ADS-B data, we resampled all trajectories at a uniform 6-second interval using the piecewise cubic Hermite interpolating polynomial (PCHIP) [29]. The trajectory dataset was then divided into training $\mathcal{D}_{\text{train}}$, validation $\mathcal{D}_{\text{val}}$, and test $\mathcal{D}_{\text{test}}$ sets using an 8:1:1 ratio (the first 80% of samples was used for training, the next 10% was used for validation, and the final 10% was used for testing).
For each data split, we constructed a past air traffic scene $\mathbf{X}$ and a future air traffic scene $\mathbf{Y}$ over 2-minute intervals (i.e., 20 time steps at 6-second intervals). This was done by sliding a window with one time step (6 seconds), as illustrated in Figure 5. To normalize the data, we applied min-max normalization to all variates using the minimum and maximum values derived from the training set. The normalization was independently applied for each variate to account for differences in their scales. As a result, we obtained a total of 509,389 traffic scenes for our experiments.
Figure 5: Illustration of air traffic scene construction using a sliding time window.
III-B Model and Training Configurations
MAIFormer consists of three identical stacked encoder layers. We set the latent dimension $D$ of the model to 256 with the FFN dimension of $4\cdot D$ equating to 1024. We configured the number of attention heads $H$ to 4 in both the MMA and AA modules. The MLP decoder consists of four linear layers with sizes of 256, 128, 64, and 32, followed by a final output size of 20 (output sequence length $S$). GELU activation functions are interspersed between the linear layers, which progressively reduce the dimensionality of the latent representation to achieve the desired sequence length. The number of variates $F$ for each agent was set to 3 (latitude, longitude, and altitude).
We trained MAIFormer using the Adam optimizer [30] over 300 epochs with a mini-batch size of 256. We set the initial learning rate to $10^{-4}$ and reduced it by a factor of 0.1 every 50 epochs. During training, we optimized MAIFormer using the L2 loss function (i.e., mean squared error), which computes the squared differences between predicted and actual trajectories. To mitigate overfitting, we employed early stopping and terminated the training loop if the validation error did not improve for 10 consecutive epochs. We developed and implemented MAIFormer using Python 3.9.16 and PyTorch 2.0.1 [31], while GPU acceleration was provided by CUDA 11.7. All experiments were conducted on a desktop with an Intel Core i7 processor, 32 GB of RAM, and a GeForce RTX 3060 GPU.
III-C Comparison Baselines
We compared the downstream performance of MAIFormer for flight trajectory prediction with that of Transformer [27] and AgentFormer [23]. To ensure fair comparison, all models are configured with the same key hyperparameters as MAIFormer, including the latent dimension, number of layers, and feedforward network dimension. To compare the prediction accuracy, we used the mean absolute error (MAE), root mean squared error (RMSE), and mean absolute percentage error (MAPE). MAE, RMSE, and MAPE were applied separately to the latitude, longitude, and altitude dimensions. MAE provides the average magnitude of absolute prediction errors, while RMSE squares the errors before averaging and gives more weight to larger errors, which makes it sensitive to outliers. MAPE was used to evaluate relative error sizes across dimensions with different scales. These evaluation metrics are defined as follows:
MAE $=\frac{1}{U}\frac{1}{N_{u}}\frac{1}{S}\sum_{u=1}^{U}\sum_{i=1}^{N_{u}}\sum_{j=1}^{S}\left|\mathbf{Y}_{u,i}^{j}-\widehat{\mathbf{Y}}_{u,i}^{j}\right|$ (13)
RMSE $=\sqrt{\frac{1}{U}\frac{1}{N_{u}}\frac{1}{S}\sum_{u=1}^{U}\sum_{i=1}^{N_{u}}\sum_{j=1}^{S}\left(\mathbf{Y}_{u,i}^{j}-\widehat{\mathbf{Y}}_{u,i}^{j}\right)^{2}}$
[…]
where $U$ is the total number of air traffic scenes in the test dataset. For each traffic scene $u\in(1,2,\ldots,U)$, $N_{u}$ is the number of agents present in the scene, and $S$ is the prediction horizon, which we set to 20. $\mathbf{Y}_{u,i}^{j}$ is the ground truth position of the $i$-th agent at the $j$-th future time step in the $u$-th scene, and $\widehat{\mathbf{Y}}_{u,i}^{j}$ is the corresponding predicted position.
TABLE I: Flight trajectory prediction performance evaluation. The best performance is highlighted in bold, and the second-best performance is underlined.
Model Horizon MAE $\downarrow$ RMSE $\downarrow$ MAPE (%) $\downarrow$
Lat () Lon () Alt (ft) Lat () Lon () Alt (ft) Lat Lon Alt
Transformer [27] 1 0.0005 0.0009 27.8757 0.0006 0.0009 37.8157 0.0014 0.0007 0.3842
5 0.0012 0.0017 71.1804 0.0013 0.0022 83.5419 0.0035 0.0016 0.9990
10 0.0025 0.0032 149.3321 0.0031 0.0039 180.3056 0.0067 0.0025 2.1641
15 0.0045 0.0054 234.2570 0.0057 0.0067 285.3934 0.0120 0.0042 3.5570
20 0.0069 0.0080 319.5657 0.0089 0.0102 389.8004 0.0187 0.0063 5.1394
AgentFormer [23] 1 0.0008 0.0015 41.6010 0.0008 0.0015 42.3061 0.0023 0.0012 0.5828
5 0.0024 0.0031 139.2278 0.0030 0.0037 168.2365 0.0066 0.0024 2.0389
10 0.0033 0.0041 187.4114 0.0044 0.0052 237.8679 0.0090 0.0032 2.8028
15 0.0034 0.0042 193.7914 0.0048 0.0057 257.6581 0.0092 0.0033 2.8636
20 0.0039 0.0054 219.3697 0.0056 0.0065 292.7384 0.0107 0.0043 3.2515
MAIFormer 1 0.0008 0.0014 29.1527 0.0008 0.0016 33.5542 0.0023 0.0011 0.4104
5 0.0012 0.0020 60.1332 0.0014 0.0021 67.7342 0.0033 0.0015 0.8347
10 0.0022 0.0030 99.7254 0.0023 0.0034 115.0246 0.0056 0.0025 1.4943
15 0.0028 0.0041 131.7546 0.0033 0.0048 159.4441 0.0078 0.0034 2.1112
20 0.0037 0.0047 164.1826 0.0046 0.0066 198.2483 0.0101 0.0037 2.7827
PI (%) - 22.46 16.93 37.93 33.33 18.14 41.23 23.01 15.27 33.85
IV Numerical Experiments
IV-A Quantitative Analysis
Table I shows the flight trajectory prediction performance of MAIFormer and all baseline models. To enable more comprehensive analysis of multi-horizon flight trajectory prediction, we report prediction errors across time horizons of 1, 5, 10, 15, and 20 steps, which correspond to 6 seconds, 30 seconds, 1 minute, 1.5 minutes, and 2 minutes into the future. We also computed the performance improvement (PI) ratio as $\left|p_{2}-p_{1}\right|/p_{2}$ to compare the best performance $p_{1}$ to the second-best performance $p_{2}$. We computed the PI ratios based on the average across all time steps of each model.
MAIFormer substantially outperformed all baseline models. The average MAE of MAIFormer shows the reductions of 22.46%, 16.93%, and 37.93% compared to the second-best model in the latitude, longitude, and altitude dimensions, respectively. Improvements in RMSE were even more pronounced, with error reductions of 33.33%, 18.14%, and 41.23% in the corresponding dimensions. We attribute these stronger RMSE gains to MAIFormer’s ability to more effectively capture the underlying dynamics of air traffic scenes. MAIFormer also achieved MAPE reductions of 23.01%, 15.27%, and 33.85% in latitude, longitude, and altitude dimensions, respectively.
Figure 6: Flight trajectory prediction results for two illustrative flight scenarios. Each row represents a different scenario, and each column shows the predictions from a different model.
When analyzing errors across different prediction horizons, the Transformer performed slightly better at early time steps (e.g., horizons 1 and 5) compared to AgentFormer and MAIFormer. The reason is likely that short-term future flight trajectories are primarily governed by the dynamics of individual aircraft, which makes them easier to extrapolate without the need to explicitly model agent interactions. In contrast, longer-term future trajectories (e.g., horizons 10, 15, and 20) are more influenced by surrounding aircraft and overall traffic patterns. Therefore, Transformer underperforms in longer-term predictions due to its inability to account for the context of neighboring traffics.
Figure 7: Visualization of attention scores for the Runway 33 arrival scenario in Figure 6. The color bar represents the magnitude of attention. Red indicates higher attention scores, and blue indicates lower scores.
IV-B Qualitative Analysis
For qualitative evaluation, we visualized the flight trajectory prediction results for two illustrative scenarios in Figure 6, which correspond to arrivals at ICN Runway 15, and Runway 33. In each plot, the blue line represents the past trajectory (used as input to the models), the green line indicates the ground-truth future trajectory, and the red line represents the model’s prediction.
For the Runway 15 arrival scenario (top row), MAIFormer accurately predicts the future trajectories of AC1, AC2, and AC3, even in the dense airspace near the runway. Both the Transformer and AgentFormer produce relatively accurate predictions for aircraft located farther from the runway (e.g., AC9 and AC10), but they struggle to maintain accuracy when predicting the trajectories of closely spaced aircraft near the runway.
For the Runway 33 arrival scenario (bottom row), both AgentFormer and MAIFormer demonstrate accurate overall predictions. However, AgentFormer fails to correctly capture the AC9’s future trajectory and incorrectly predicts that it will arrive ahead of AC8, which is already established on a structured arrival flow. In contrast, MAIFormer successfully captures the subtle interactions between the two aircraft and correctly places AC9 behind AC8 in the sequence. This example highlights MAIFormer’s better capability of understanding traffic scenes and inter-agent interaction modeling.
Figure 7 illustrates the attention distribution across different aircraft computed by AgentFormer and MAIFormer for the Runway 33 arrival scenario (Figure 6, bottom row). The attention scores were visualized from the perspective of AC9 (which serves as the query), and the color reflects how much attention it assigns to each surrounding aircraft. AgentFormer computes attention scores at the trajectory point level and models the relationship between individual points across different time steps. As a result, the attention score can vary even within the same aircraft’s trajectory. This results in fine-grained information, but the excessively detailed attention distributions are inherently difficult for human operators to interpret. As the number of aircraft in a traffic scene increases, the complexity of interpreting these point-to-point relationships grows rapidly, which reduces their practical utility in operational contexts.
In contrast, MAIFormer computes attention at the agent level and assigns a single attention score to each aircraft. These agent-to-agent attention values are far more interpretable as they represent the relative importance of each aircraft as perceived by the querying aircraft (AC9 in this case). This abstraction enables a more intuitive and human-understandable interpretation of inter-agent relationships, which is particularly valuable in safety-critical domains such as ATM. Interestingly, MAIFormer assigns high attention not only to AC5 and AC8 but also to AC4. A possible explanation for this behavior is that AC9 may have had the potential to enter the arrival sequence earlier before AC5, depending on the location and speed of AC4. The model appears to have incorporated about this possibility and thus assigned a relatively high attention weight to AC4.
V Conclusion
We have proposed MAIFormer as a multi-agent flight-trajectory prediction model that learns both the spatio-temporal patterns of individual aircraft and the interaction patterns between them. Its effectiveness was demonstrated using real-world air traffic surveillance data, and the experimental results showed that it achieves better prediction accuracy compared to baseline models. Its hierarchical attention architecture allows it to produce interpretable attention distributions that explicitly indicate the degree of influence that each surrounding aircraft has on a given aircraft’s future trajectory.
Nevertheless, despite its strong performance, several aspects require further exploration to improve its practical applicability in real-world air traffic operations. While MAIFormer provides insights into aircraft interactions through its agent attention, it remains uncertain whether these learned patterns truly reflect a human ATC’s attention. Future research should involve human-in-the-loop simulations or expert assessments to evaluate the interpretability and operational alignment of the attention distributions produced by the proposed method. Furthermore, the model relies solely on flight trajectory data, but in real-world operations, external contextual factors such as meteorological conditions, flight procedures, and airspace constraints play a significant role in ATC decision making and aircraft behavior. Incorporating such conditional information into the model would improve prediction accuracy and bring the model closer to the standard required for real-world deployment in an ATM system.